OpenStudy (anonymous):

find tangent line x^3+y^3=y+21 at (3,-2) I got y=-27/11x +59/11 but it's saying its not right, and I cant see the error in my ways

OpenStudy (anonymous):

oh hmm shouldnt it be 3x^2 +3y^2(dy/dx)=dy/dx and after solving that its the slope?

myininaya (myininaya):

yep! :)

myininaya (myininaya):

\[3y^2 y'-y'=-3x^2\]

myininaya (myininaya):

\[y'(3x^2-1)=-3x^2\]

myininaya (myininaya):

Now divide both sides by \[3x^2-1\]

OpenStudy (zarkon):

\[3y^2-1\]

myininaya (myininaya):

oops type-0

myininaya (myininaya):

yes y doesn't usually change to x unless some kind of act of god happens

myininaya (myininaya):

or myininaya screws up

OpenStudy (zarkon):

I like doing it the quick way \[\frac{dy}{dx}=-\frac{\frac{\partial f}{\partial x}}{\frac{\partial f}{\partial y}}\]

OpenStudy (anonymous):

wait what you just typed to me is the slope? cuz thats dy/dx implicit d confuses the hell out of me

myininaya (myininaya):

just use y' instead

OpenStudy (anonymous):

dy/dx is the slope?

OpenStudy (anonymous):

y'

myininaya (myininaya):

y' always looks prettier to me

myininaya (myininaya):

\[y'=\frac{dy}{dx}\]

OpenStudy (anonymous):

cuz the equation I got up top is the tangent line with the slope as y'

myininaya (myininaya):

wait what is your question? i'm confused

OpenStudy (anonymous):

hahaha okay so how to do I get the slope of the tangent line?

myininaya (myininaya):

\[slope=y'=\frac{dy}{dx}=\frac{-3x^2}{3y^2-1}\]

OpenStudy (anonymous):

hmmm thats -27/11 and it isnt giving me the correct answer

OpenStudy (anonymous):

the tangent line I posted is it with that slope so idk what i did wrong

myininaya (myininaya):

\[y'|_{(3,-2)}=\frac{-3(3)^2}{3(-2)^2-1}=\frac{-27}{11}\] I don't think I understand your question yes that is the right slope

myininaya (myininaya):

\[y=\frac{-27}{11}x+b\] What point do we know that is on this line (3,-2) plug this in to find b

OpenStudy (anonymous):

okay so if that is my slope then y+2=-27/11(x-3) y= -27/11x +59/11 but it is saying that is incorrect

OpenStudy (ash2326):

@Awstinf try typing it in standard form \[27x+11y=\frac{59}{11}\]

OpenStudy (ash2326):

Sorry\[27x+11y=59\]

myininaya (myininaya):

That is right lol

myininaya (myininaya):

Ok well that just means you should subtract 2 both sides

myininaya (myininaya):

\[y=\frac{-27}{11}(x-3)-2\] but you already did

myininaya (myininaya):

so it just says y=?

OpenStudy (anonymous):

yeah :\

myininaya (myininaya):

So it does it say to put in slope intercept form?

OpenStudy (anonymous):

yeah, but webwork is kind of screwy. Probably one of the worst systems I've ever dealt with. Wolfram should make something like it so we never have to deal with it ever again lol.

myininaya (myininaya):

I'm afraid I don't know what to tell you Your answer is right

OpenStudy (anonymous):

thank you for helping me! again!